Operating Systems and Systems Integration

Setting up Netscape

1 Aim

Here we set up Netscape so that when we click on a hyperlink to a pdf (“Acrobat”) file, the program xpdf will start up to read it. We also configure the proxy configuration, set the font size bigger, and make http://nicku.org/ossi/ the default home page.

2 Background

My dictionary says a proxy is an agent given responsibility to act for another. A proxy server is a network server that sits on the border of the network. It works as a “gobetween” between a web browser (for example, in the Campus), and a web server, out on the Internet. It works like this:

• request from web browser goes to proxy

• If proxy has a recent copy of the web page in its cache, it sends the page straight to the web browser,

• If the proxy has no copy,

∘ It asks the web server for a copy

∘ It stores the page in its cache

∘ It sends the page to the web browser.

The proxy server has two purposes:

1. It speeds up access by web clients to popular web pages by caching them locally

2. It increases security, since the gateway to the Internet can be better controlled; the proxy is a component in the firewall, which filters dangerous traffic to reduce the risk of bad crackers out on the Internet breaking into important computers in the College.

3 Procedure

3.1 Setting up Netscape to start xpdf whenever you click on a hyperlink to a PDF file

Note: I suggest that you use the acrobat reader instead. The instructions for this are

given at http://intra.tyict.vtc.edu.hk/ossi/lab/install-acrobat/install-acrobat.pdf.

Please do that instead.

Setting up Netscape Operating Systems and Systems Integration

2

1. Start Linux

2. start Netscape by clicking once on the Netscape icon in the panel at the bottom of the screen.

3. From the Edit menu select Preferences, then click on that ▶ triangle next to Navigator to turn it into a ▼ triangle, and select Applications.

4. Scroll through the list of applications until you see the Description Portable Document Format. Select that line, and click on the Edit… button.

5. Click on the Application button.

6. Enter /usr/bin/xpdf %s in the Application field.

7. The screen will look something like figure 1.

Figure 1: Setting up Netscape to run xpdf when you click on a hyperlink to a PDF file.

3.2 Entering the proxy setting

Unless the proxy is set, you will only be able to connect to other web servers behind the College firewall. To set it:

1. From the Edit menu select Preferences, then click on that ▶ triangle next to Advanced to turn it into a ▼ triangle, and select Proxies.

2. Click on the Automatic Proxy Configuration button, and in the field there type

   http://tyintra.vtc.edu.hk/proxy.proxy, as in figure 2 on the next page.

This will download a Javascript file (you will find it later, after you use Netscape, in the file ∼/.netscape/proxyconf. The Javascript selects a suitable proxy depending on what destination you want to connect to, and where you are.

3. Click on the Reload button.

4. Click on OK.

Setting up Netscape Operating Systems and Systems Integration

3

Figure 2: The automatic proxy configuration in Netscape.

3.3 Setting the Caching policy

The web browser stores web pages that you download on the hard disk in the web cache. When you visit the page a second time, you will see the copy rather than wait for the browser to download the page a second time. This speeds things up. There is a problem, however, when the content changes between the first time you read the page to the second time. Normally you will not see the changed version, only the first version.

The teaching material at our subject web site changes often. To make sure that you are getting the latest version, please do the following:

Figure 3: The cache setting in Netscape. This is also important for Web application developers. For people like us (system administrators and web site developers), this is nearly always the right choice.

(a) From the Edit menu select Preferences.

Setting up Netscape Operating Systems and Systems Integration

4

(b) click on the ▶ triangle next to Advanced to turn it into a ▼ triangle, and select Cache.

(c) Select “Every time, instead of “Once per session”, as in figure 3 on the preceding page.

(d) Click on the OK button. Note that there is a slight performance penalty, but it is worth it to reduce the confusion with rapidly changing web sites, and for those who are developing them.

3.4 Setting home page, larger font

There are two other settings that are convenient for us using Netscape here; one is to select the home page for our own web server:

(a) From the Edit menu select Preferences, then click on Navigator.

(b) For the Home page Location, enter: http://nicku.org/ossi/

(c) To make the default font larger, click on the ▶ triangle next to Appearance to turn it into a ▼ triangle, and select Fonts.

(d) Change the Size to 18.

(e) Click OK, and you are ready to use Netscape to read the teaching material on our web site.